The distance of separation of two spheres of masses m and 5m and radius r and 2r in space is 12r if they attract each other due to gravitational forces then the distance coveres by the smaller sphere before collison is?

Grade:12

1 Answers

Arun
5 years ago

 

Since there is no external force acting on the system of two masses M and 5M. Thus the centre of mass of the system will remain unaltered.

Considering mass M to be initially at the origin, The CM will be

X = (0×M + 5M×12R)/(5M+M)

=> X = 10R

Thus CM is at a distance 10R which remains unchanged before and after collision.

At collision the centre of the spherical masses will be 3R apart from each other

Now let the collision take place at a distance x from the origin

 

After collision

=> {mx + 5m(x+3R)}/6m = 10R

=> 6x + 15R = 60R

=> 6x = 45R

=> x = 7.5R